OpenStudy (anonymous):

Graph the line with slope 3 passing through the point −3,-5

OpenStudy (anonymous):

its -3 and -5 not a postive 3

OpenStudy (campbell_st):

well the line, in slope intercept form is y = mx + b and you are told m = 3 so substitute x = -3 and y = -5 to find b when you get the line, use a table of values such as x: -1 : 0 : 1 substitute each x value to find the matching y. ------------- x = -1 y = 3*(-1) + 4 so y = 1 y : 1 : then plot the 3 points hope it helps